OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN Given: F(x) = 3x^2 + 1, G(x) = 2x - 3, H(x) = x F(3) + G(4) - 2H(5) = A: 13 B: 23 C: 33

OpenStudy (anonymous):

well that's easy , it's clearly b

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

go to www.wolframalpha.com and it will show you the answer

OpenStudy (anonymous):

well you see , look at the given formula , quack , your question is irrelevant

OpenStudy (anonymous):

fan and medal thank you

OpenStudy (agent0smith):

F(x) = 3x^2 + 1, G(x) = 2x - 3, H(x) = x F(3) + G(4) - 2H(5) = first find F(3), by plugging in x=3 into F(x) = 3x^2 + 1 then find F(3), by plugging in x=4 into G(x) = 2x - 3 then find H(5), by plugging in x=5 into H(x) = x Then use those results to find F(3) + G(4) - 2H(5)

OpenStudy (anonymous):

^ that's what i said